Two angles are supplementary and the measure of one of them is 20% less than three times the measure of the other. Find the measure of each angle.
Found 2 solutions by checkley71, richmusick:
Answer by checkley71(8403)

X+Y=180 OR X=180-Y
X=.2(3Y)
180-Y=.6Y
1.6Y=180
Y=180/1.6
Y=112.5 FOR THE LARGER ANGLE.
X=.2(3*112.5)
X=.2(337.5)
X=67.5 FOR THE SMALLER ANGLE.
PROOF
112.5=67.5=180
180=180

Answer by richmusick(3)

X+Y=180 OR X=180-Y
X=3Y - .2(3Y)
X=3Y - .6Y
X=2.4Y
2.4Y + Y = 180
3.4Y=180
Y=180/3.4
Y=53° FOR THE SMALLER ANGLE.
X+53°=180°
X=180° - 53°
X=127° FOR THE LARGER ANGLE.
PROOF
Y = 53°
X = 2.4Y
2.4(53) = 127°
53° + 127°=180°